OpenStudy (anonymous):

Add or subract and simplify the answer need the step please. thanks in advance 5/(6a^2 )+ 1/(6ab)

OpenStudy (anonymous):

5/(6a^2 )+ 1/(6ab) need the same denominator, = 5/(6a^2 )+ a/(6a^2b) =(5+a)/(6a^2b)

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

5B+a/(6a^2b)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

always welcomed